# Language Models can Self-Improve at State-Value Estimation for Better Search

### v1

@@ Abstract
Collecting ground truth task completion rewards or human demonstrations for multi-step reasoning tasks is often cost-prohibitive and time-consuming, especially in interactive domains like web tasks. To address this bottleneck, we present self-taught lookahead, a self-supervised method that leverages state-transition dynamics to train a value model capable of effectively guiding language model-controlled search. We find that moderately sized (8 billion parameters) open-weight value models improved with self-taught lookahead can match the performance of using a frontier LLM such as gpt-4o as the value model. Furthermore, we find that self-taught lookahead improves performance by 20% while reducing costs 37× compared to previous LLM-based tree search, without relying on ground truth rewards.
Machine Learning, ICML
@@ 1 Introduction
While current large language models (LLMs) have exhibited impressive intrinsic reasoning capabilities (Wei et al., 2022; Wang et al., 2022), recent work (Yao et al., 2024; Xie et al., 2024) has demonstrated that eliciting more deliberate reasoning through explicit tree search improves performance on complex multi-step reasoning tasks. During this search process, a policy LLM proposes possible actions from the current state, and a value LLM evaluates and selects the most promising states to explore. Recent work has found the strength of the value model to be strongly correlated with overall performance (Chen et al., 2024; Liu et al., 2024a).
Much prior work on LLM-based tree search simply prompts an off-the-shelf LLM to serve as both the policy and value models during the search process, where performance is constrained by the capabilities of this model. Recent approaches have addressed this constraint by utilizing ground truth task completion rewards during Monte Carlo tree search (MCTS) to guide node selection. Zhou et al. (2024) show that MCTS using these rewards can outperform gradient-based methods that train on human demonstrations or rejection sampled trajectories. However, collecting ground-truth rewards can be costly. For example, in interactive domains like web tasks, data collection can cost thousands of dollars (Yao et al., 2022a). This constraint motivates the need for new methods that can perform self-supervised training of the policy or value LLM without ground truth rewards.
In this paper, we introduce self-taught lookahead (STL), a technique that enables an LLM-based value model to improve itself by bootstrapping an initial value function and leveraging the environment’s state transition dynamics without requiring any ground-truth rewards. To the best of our knowledge, this is the first demonstration that an LLM-based value function can self-improve without labels or rewards. We show that using this self-improved value model during search can outperform computationally expensive MCTS methods (Zhou et al., 2024), while reducing inference costs by an order of magnitude.
@@
The self-taught lookahead (STL) process (Figure 1) begins by generating self-improvement data through a single step of lookahead within a tree search. Analogous to the Bellman update, this lookahead refines the estimated value of a state by leveraging information about potential future states. However, unlike classical reinforcement learning (RL) methods, which rely on explicit environment rewards, STL uses a large language model (LLM) to estimate state values. Specifically, we fine-tune a value model to reason about the utility of a state by predicting the next best action, its resulting state, and a corresponding rationale for the value of that state. During training, the model is fine-tuned using rollouts of states and actions within the environment. At inference time, the value model simulates a step of lookahead to provide more accurate value judgments. STL, therefore, more explicitly captures the paradigm of “verbal reinforcement learning” than previous work (Shinn et al., 2024) by learning from a natural language representation of the mechanics of a traditional RL algorithm.
@@
Figure 1: Self-taught lookahead self-improves the value model by learning from state-transition dynamics. During the data generation phase (top left), tree search is used to discover diverse states. For every observed state $s$ encountered during the search, successor states are expanded using base policy $\pi_{\theta}$ and the current value model $V_{\phi_{k}}$, and a formatted textual training example is formed using verbal representations of the next best action and successor state, as well as $V_{\phi_{k}}$’s outputted value reasoning ($r$) and numerical value ($v$) (top middle). These examples are used to fine-tune $V_{\phi_{k+1}}$, which will be used in the next iteration of the algorithm (top right). Value models learned during self-taught lookahead can be used to evaluate unseen states encountered during search on unseen tasks by simulating a step of lookahead including the next best action and the best successor state $\tilde{s}^{\prime}$ (bottom).
By learning to simulate the entire process of lookahead rather than only iterated values, STL can leverage the strong generalization improvements of LLMs resulting from this sort of rationalization (Nye et al., 2021; Zelikman et al., 2022). Our results (§4) demonstrate that tree search with an STL fine-tuned llama-3.1-8b-instruct value model improves performance by $39\%$ or more compared to the base model and matches the performance of search with a base gpt-4o value model on unseen tasks. We show that these results hold across web agent and math reasoning domains, including WebShop (Yao et al., 2022a) and Game-of-24 (Yao et al., 2024).
Finally, in §5, we demonstrate that search with STL value models achieves Pareto optimality when considering cost and environment usage (states expanded). The results reveal that self-taught lookahead effectively transfers computation from expensive large closed-source models to cheaper open-source alternatives, achieving a 37$\times$ cost reduction while outperforming previous LLM tree search methods (Zhou et al., 2024) by 20%. Additionally, we analyze scaling trends for self-taught lookahead and find that STL can even be utilized to improve smaller models ($\leq$ 3 billion parameters).
2 Background
2.1 Guiding Tree Search with Language Models
@@
Within a state space $\mathcal{S}$, the goal of the tree search is to reach a desired state $s_{*}\in\mathcal{S}$ from an initial state $s_{0}\in S$, where $s_{*}$ is determined based on the natural language task $x\in\mathcal{T}$. A state $s_{i}\in S$ might be a step in a reasoning chain or an intermediate webpage in a web navigation task. While the algorithmic details vary based on the tree search method e.g. breadth-first search (BFS) or Monte Carlo Tree Search (MCTS), to adapt these methods to utilize language models (LMs), we simply need to define how new states (successors) are generated and evaluated.
@@
$a^{(j)}_{i}\sim\pi_{\theta}(a_{i}|x,s_{0},\cdots,s_{i}),\forall j\in\{1,B\}$ (1)
@@
A value $v_{s_{i}|x}$ and a rationale for the value $r_{s_{i}|x}$ of a state $s_{i}$ is generated using a LLM-based value model $V_{\phi}:\mathcal{T}\times\mathcal{S}^{*}\rightarrow\mathbb{R}\times\mathcal{L}$, where $\mathcal{L}$ is the space of natural language sequences:
$(r_{s_{i}|x},v_{s_{i}|x})\sim V_{\phi}(x,s_{0},\cdots,s_{i})$ (2)
Note that because $V_{\phi}$ is an LLM, $r_{s_{i}|x}$ is generated first, and $v_{s_{i}|x}$ is then conditionally generated. We will denote $r_{s_{i}|x}\sim\text{Res}(V_{\phi}(x,s_{0},\cdots,s_{i}))$ and $v_{s_{i}|x}\sim\text{Val}(V_{\phi}(x,s_{0},\cdots,s_{i}))$. While previous work (Zhou et al., 2024; Koh et al., 2024b; Yao et al., 2024; Yu et al., 2024) do generate rationales during state evaluation, usually they are generated purely to leverage the performance improvements of LLMs when asked to rationalize (Wei et al., 2022) and are subsequently discarded. As described in §3, we explicitly use these rationales for self-improvement.
@@ 3 Better State-Value Estimation with Self-Taught Lookahead
Unlike neural models traditionally used for state-value estimation in the learning to search literature (Silver et al., 2017), an LLM can leverage both conventional numerical values and natural language reasoning to estimate state value. Through our STL method, we train LLM value models that learn to better assign values to states based on their expected future utility by constructing and learning from rationales that explicitly capture domain-specific state transition dynamics. For instance, without understanding these transitions, it might not be clear whether a Close (X) button on a website interface exits the current view or the entire workflow [^1] This is a known problem in UX design. in web tasks . While MCTS-based methods (Zhang et al., 2024a) may learn how transitions affect value assignment implicitly i.e. through the backup of terminal state rewards, we explicitly learn from these transitions via supervised fine-tuning (SFT) on natural language representations of future actions and states, and value estimation rationales. We find that this explicit approach allows us to employ more compute-efficient search methods (§5) while maintaining or improving performance (§4) and leverage the demonstrated generalization improvements of learning from rationales (Nye et al., 2021; Zelikman et al., 2022).
Learning better state-value estimates from the state transition dynamics is also ideal for self-improvement as the environment provides state transition outcomes, so our approach does requires neither ground truth reward, human demonstrations, nor labels. STL thus assumes a static policy model $\pi_{\theta}$ and only trains the value model through one or more iterations of self-improvement. We denote the value model initialized with a base LLM $V_{{\phi}_{0}}$ and the value model used in a subsequent iteration $k$ to assign values and generate rationales $V_{\phi_{k}}$. Figure 1 presents the method in full, but this section enumerates the data generation, training, and inference steps in more detail (see Algorithm 1 in Appendix A for a concisely written algorithm).
@@ 3.1 Generating Rollouts
An iteration $k$ of self-taught lookahead starts with a dataset $\mathcal{D}_{\texttt{rollout}_{k}}$ of natural language tasks for the current iteration. For each $x_{i}\in\mathcal{D}_{\texttt{rollout}_{k}}$, we roll out the search tree using $\pi_{\theta}$ and $V_{\phi_{k}}$. Using tree search enables us to collect a diverse set of states so that the value model trained on these states’ values can better generalize to unseen states and tasks. We empirically demonstrate this generalization in §4. When visiting state $s_{j}$ on the trajectory $\{s_{0},\cdots,s_{j}\}$ during tree search, we compute $s_{j}$’s lookahead value, $y_{s_{j}}$:
$y_{s_{j}}\leftarrow\gamma\max_{a\in A_{s_{j}}}\Big{\{}\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\cdots,s_{j},s_{j+1}))\Big{\}}$ (3)
where $\gamma$ is the discount factor and $s_{j+1}\leftarrow T(s_{j},a)$. These lookahead values capture a better estimate of the true value of $s_{j}$ as they account for $s_{j}$’s successor states. In §6, we describe how generating and learning from these $y$’s is similar to fitted value iteration.
@@
To better capture the state transition dynamics, we also generate action-outcome rationales when visiting a state $s_{j}$. These rationales are of the form “{action} {outcome_state} {value_rationale}” where action is sampled from the optimal actions from the $\max$ in (3), outcome_state is the successor state derived from taking the action, and value_rationale is the rationale for the evaluation of this successor state generated by $V_{\phi_{k}}$. Fine-tuning on these rationales will enable a value model to predict the result of taking an action and incorporate this prediction (lookahead) into the current state’s value estimate. Formally, we can define these action-outcome rationales $o_{s_{j}}$:
$o_{s_{j}}\leftarrow a^{*}_{j}||s^{*}_{j+1}||\text{Res}(V_{\phi_{k}}(x_{i},s_{0},\cdots,s_{j},s_{j+1}))$ (4)
@@
$a^{*}_{j}\in\operatorname*{arg\,max}_{a\in A_{s_{j}}}\Big{\{}\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\cdots,s_{j},T(s_{j},a)))\Big{\}}\$ (5)
The training data set for iteration $k$ is thus a set of tuples: $\mathcal{D}_{k}=(s_{k},o_{s_{k}},y_{s_{k}})$. Depending on the task, it might be necessary to automatically filter out tuples that have malformed rationales or account for the same state seen multiple times in different iterations (see §4.2).
@@ 3.2 Fine-Tuning the Value Model
We start training the new value model $V_{\phi_{k+1}}$ from the initial or base LLM value model $V_{\phi_{0}}$. We can then train using standard supervised fine-tuning negative log-likelihood loss for the generation of both the action-outcome rationale and the lookahead value ($o_{s}||y_{s}$) of the state. We train the value model to generate the rationale before estimating the value. Automatically constructed text or formatting as seen in Figure 1 are applied for easier learning.
@@ 3.3 Search after Self-Taught Lookahead
A value model resulting from iteration $k$ of self-taught lookahead ($V_{\phi_{k+1}}$) can directly replace a value model in any search algorithm: §4.2 and §4.1 demonstrates usage in BFS and Greedy search, respectively. As shown in Figure 1 (bottom), $V_{\phi_{k+1}}$ simulates a step of lookahead for the state $s_{n}$ i.e. for $({r}_{s_{n}|x},{v}_{s_{n}|x})\sim V_{\phi_{k+1}}(x,s_{0},\cdots,s_{n})$,
@@
where $\tilde{a}_{{n+1}}||\tilde{s}_{{n+1}}$ is a simulated step of lookahead and $\tilde{r}_{s_{n+1}|x}$ is the value rationale of the simulated successor.
@@ 4 Experiments
We benchmark the self-taught lookahead self-improvement approach on applied web agent tasks and traditional math reasoning tasks [^2] Our code is available at https://github.com/ethanm88/self-taught-lookahead..
Table 1: Comparison of methods for the WebShop task. For each method, the models and inference configuration are listed along with the usage of human demonstrations, ground truth rewards, environmental observations, and fine-tuning marked with ✓ (used) and ✗ (unused). In the Reward and Demo Free setting, only environment observations are accessible.
Method Inference Config Human Demo. GT Reward Env. Obs. FT
Reflexion (Shinn et al., 2024)
LATS (Zhou et al., 2024) MCTS
IL (Yao et al., 2022a) pass@$3$
IL+RL (Yao et al., 2022a) pass@$3$
MCTS + DPO (Putta et al., 2024) MCTS
Greedy Baseline Greedy (pass@$3$)
MCTS Baseline MCTS (pass@$3$)
STL (Ours) Greedy (pass@$3$)
Expert (Yao et al., 2022a)

### v2

@@ Abstract
Collecting ground-truth rewards or human demonstrations for multi-step reasoning tasks is often prohibitively expensive and time consuming, especially in interactive domains like web tasks. To address this bottleneck, we present self-taught lookahead (STL), a self-supervised method that leverages state-transition dynamics to improve a value model capable of effectively guiding language model-controlled search without any labeled data. We find that moderately sized (8 billion parameters) open-weight value models improved with STL can match the performance of using a gpt-4o value model. Furthermore, we find that specialized value models learned with STL can be deployed with computationally lightweight search algorithms, achieving performance that matches that of more expensive tree search methods, while reducing costs by an order of magnitude.
@@ 1 Introduction
While large language models (LLMs) demonstrate strong reasoning capabilities by generating extended token sequences before answering wei2022chain ; guo2025deepseek ; muennighoff2025s1 , guiding inference with explicit tree search has the potential to further improve performance on tasks with a structured state space yao2024tree ; xie2024self . In this setting, a policy LLM proposes candidate actions, and a value LLM evaluates resulting states to steer the search toward promising trajectories. Figure 1 summarizes the assumptions different LLM-driven search methods make about information available during training and inference. Reward and Demo Learning methods yao2022webshop ; zhang2024rest_mcts ; jin2025search ; song2025r1 assume access to ground-truth reward signals or human demonstrations and optimize the model with reinforcement learning (RL) or imitation learning (IL). In contrast, Reward-Guided Inference strategies zhou2023lats ; shinn2024reflexion forego explicit reward during training; they only consult a reward signal at inference time, using it to guide procedures such as LLM-based Monte-Carlo Tree Search (MCTS). However, collecting ground-truth rewards or human demonstrations may not be possible in every environment, and can oftentimes be costly. For example, for web agent tasks, even small-scale data collection can cost thousands of dollars yao2022webshop . Reward and Demo Free methods yao2022react ; yao2024tree relax this assumption as they can operate without access to reward. However, these methods often rely on prompting an off-the-shelf LLM to serve as both the policy and value models during the search process, which constrains performance to the capabilities of the underlying LLM.
In this paper, we introduce Self-Taught Lookahead (STL), a Reward and Demo Free, self-supervised framework for interactive, multi-step reasoning tasks. Building on evidence that search quality is strongly influenced by the accuracy of the state-value estimator (chen2024search_useful, ; liu2024don, ), STL improves an LLM-based value function for better search performance. Unlike neural models traditionally used for state-value estimation in the learning to search literature silver2017masteringall , an LLM can leverage both conventional numerical values and natural language reasoning to estimate state values. STL exploits this feature by having an LLM value model learn to better assign values to states based on their expected future utility by constructing and learning from rationales that explicitly capture state transition dynamics. For instance, without understanding these transitions, it might not be clear whether a Close (X) button on a website interface exits the current view or the entire workflow in web tasks harley2019cancel . Learning better state-value estimates from state transition dynamics is especially well-suited for self-improvement in agentic tasks, where the environment directly provides transition outcomes. As a result, our approach requires neither ground truth rewards nor human demonstrations.
@@
STL (Figure 2) begins by generating self-improvement data through a single step of lookahead within a tree search. Analogous to the Bellman update, this lookahead refines the estimated value of a state by leveraging information about potential future states. However, unlike classical reinforcement learning (RL) methods, which rely on explicit environment rewards, STL uses a large language model (LLM) to estimate state values. Specifically, during STL, a value LLM is fine-tuned to reason about the utility of a state by predicting the next best action, its resulting state, and a corresponding rationale for the value of that state. During training, the model is fine-tuned using rollouts of states and actions within the environment. At inference time, instead of taking a step of lookahead in the environment, the improved value model simulates a step of lookahead to provide more accurate value judgments.
By representing the lookahead process in natural language rather than regressing solely on value estimates, STL takes advantage of LLMs’ strong generalization to unseen tasks via learned textual reasoning nye2021show ; zelikman2022star . For instance, our results (§4) on web agent tasks from WebShop yao2022webshop demonstrate that tree search with an STL fine-tuned llama-3.1-8b-instruct value model improves performance by $39\%$ or more compared to the base llama model. Furthermore, STL matches the performance of search with a base gpt-4o value model and even achieves comparable results to Reward-Guided Inference methods such as LATS zhou2023lats on unseen tasks. We show that these results also hold for math puzzles and multi-hop question answering yao2024tree ; yang-etal-2018-hotpotqa . Finally, by enabling the use of a small open-source value model during inference, STL leads to significant cost reduction when agents are deployed. Through an efficiency analysis (§5), we find that STL costs 5$\times$ less than a similarly performing gpt-4o value model. As STL produces more accurate state-value estimates, the resulting value models can effectively guide search algorithms that expand fewer states, while maintaining strong performance.
Figure 1: The information accessible during learning and inference across common search settings, exemplified using web tasks. Our Self-Taught Lookahead method is Reward and Demo Free, yet is able to self-improve by learning from state transitions in the form of lookahead values and rationales.
@@
Figure 2: Self-taught lookahead self-improves the value model by learning from state-transition dynamics. During the data generation phase (top left), tree search is used to discover diverse states. For every observed state $s$ encountered during the search, successor states are expanded using base policy $\pi_{\theta}$ and the current value model $V_{\phi_{k}}$, and a textual training example is formed using verbal representations of the next best action and successor state, as well as $V_{\phi_{k}}$’s outputted value reasoning ($r$) and numerical value ($v$) discounted by $\gamma$ (top middle). These examples are used to fine-tune $V_{\phi_{k+1}}$, which will be used in the next iteration of the algorithm (top right). Value models learned during STL can be used to evaluate unseen states encountered during search on unseen tasks by simulating a step of lookahead, including the next best action and the best successor state $\tilde{s}^{\prime}$ (bottom).
2 Background: Guiding Tree Search with Language Models
@@
Within a state space $\mathcal{S}$, the goal of the tree search is to reach a desired state $s_{*}\in\mathcal{S}$ from an initial state $s_{0}\in S$, where $s_{*}$ is determined based on the natural language task $x$, with $x\in\mathcal{X}$, the set of all possible tasks. A state $s_{i}\in S$ might be a step in a reasoning chain or an intermediate webpage in a web navigation task. While the algorithmic details vary based on the tree search method e.g., breadth-first search (BFS) or MCTS, to adapt these methods to utilize language models (LMs), we simply need to define how new states (successors) are generated and evaluated.
@@
$a^{(j)}_{i}\sim\pi_{\theta}(a_{i}|x,s_{0},\ldots,s_{i}),\forall j\in\{1,\ldots,B\}$ (1)
@@
A value $v_{s_{i}|x}$ and a rationale for the value $r_{s_{i}|x}$ of a state $s_{i}$ is generated using a LLM-based value model $V_{\phi}:\mathcal{X}\times\mathcal{S}^{*}\rightarrow\mathbb{R}\times\mathcal{L}$, where $\mathcal{S}^{*}$ is the set of all finite sequences over the state space $S$ and $\mathcal{L}$ is the space of natural language sequences:
$(r_{s_{i}|x},v_{s_{i}|x})\sim V_{\phi}(x,s_{0},\ldots,s_{i})$ (2)
Note that because $V_{\phi}$ is an LLM, it is usually prompted in a chain-of-thought wei2022chain manner to first generate $r_{s_{i}|x}$ and then conditionally generate $v_{s_{i}|x}$ based on the rationale. For notational simplicity, will denote these two generated entities as $r_{s_{i}|x}\sim\text{Rat}(V_{\phi}(x,s_{0},\ldots,s_{i}))$ and $v_{s_{i}|x}\sim\text{Val}(V_{\phi}(x,s_{0},\ldots,s_{i}))$. While previous work zhou2023lats ; koh2024tree ; yao2024tree ; yu2024exact do generate rationales during state evaluation, usually they are generated purely to leverage the performance improvements of LLMs when asked to rationalize and are subsequently discarded. As described in §3, inspired by self-taught reasoning zelikman2022star , we explicitly use these rationales during self-improvement by fine-tuning on them along with the value to learn how to better estimate state values in a given domain.
@@ 3 Better State-Value Estimation with Self-Taught Lookahead
In this section, we present our proposed self-taught lookahead method. §3.1 and §3.2 describe the STL improvement procedure, while §3.3 explains how a STL-improved value model operates during inference. See Figure 2 and Algorithm 1 in Appendix A for an overview of the method.
@@ 3.1 Generating Rollouts
STL assumes a static policy model $\pi_{\theta}$ and only trains the value model through one or more iterations of self-improvement. We denote the value model initialized with a base LLM $V_{{\phi}_{0}}$ and the value model used in a subsequent iteration $k$ to assign values and generate rationales $V_{\phi_{k}}$.
An iteration $k$ of STL starts with a dataset $\mathcal{D}_{\texttt{rollout}_{k}}\subset\mathcal{X}$ of natural language tasks for the current iteration. For each $x_{i}\in\mathcal{D}_{\texttt{rollout}_{k}}$, we roll out the search tree using $\pi_{\theta}$ and $V_{\phi_{k}}$. Using tree search enables us to collect a diverse set of states so that the value model trained on these states’ values can better generalize to unseen states and tasks. We demonstrate this generalization in §4. When visiting state $s_{j}$ on the trajectory $\{s_{0},\ldots,s_{j}\}$ during tree search, we compute $s_{j}$’s lookahead value, $y_{s_{j}}$:
$y_{s_{j}}\leftarrow\gamma\max_{a\in A_{s_{j}}}\Big{\{}\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},T(s_{j},a)))\Big{\}}$ (3)
where $\gamma$ is the discount factor. Since tasks are episodic, we set $\gamma=1$ bernstein2001reinforcement ; van2005survey ; leffler2007efficient . These lookahead values capture a better estimate of the true value of $s_{j}$ as they account for $s_{j}$’s successor states. In §6, we describe how generating and learning from these $y$’s is similar to fitted value iteration.
@@
To better capture the state transition dynamics, we also generate action-outcome rationales when visiting a state $s_{j}$. These rationales are of the form “{action} {outcome_state} {value_rationale}” where action is selected by the $\max$ operator in Equation 3, outcome_state is the state observed after taking the action in the environment and value_rationale is the rationale for the evaluation of this successor state generated by $V_{\phi_{k}}$. Fine-tuning on these rationales will enable a value model to predict the result of taking an action and incorporate this prediction (lookahead) into the current state’s value estimate. Formally, we can define these action-outcome rationales $o_{s_{j}}$:
$o_{s_{j}}\leftarrow a^{*}_{j}||s^{*}_{j+1}||\text{Rat}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},s^{*}_{j+1}))$ (4)
@@
$a^{*}_{j}\in\operatorname*{arg\,max}_{a\in A_{s_{j}}}\Big{\{}\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},T(s_{j},a)))\Big{\}}\$ (5)
The training data set for iteration $k$ is thus a set of tuples: $\mathcal{D}_{k}=(s_{k},o_{s_{k}},y_{s_{k}})$. Depending on the task, it might be necessary to automatically filter out tuples that have malformed rationales or account for the same state seen multiple times in different iterations (see Appendices DE, and F).
@@ 3.2 Fine-Tuning the Value Model
We start training the new value model $V_{\phi_{k+1}}$ from the initial or base LLM value model $V_{\phi_{0}}$. We can then train using standard fine-tuning negative log-likelihood loss for the generation of both the action-outcome rationale and the lookahead value ($o_{s}||y_{s}$) of the state. We train the value model to generate the rationale before estimating the value. Automatically constructed text or formatting, as seen in Figure 2, is applied for easier learning.
@@ 3.3 Search after Self-Taught Lookahead
A value model resulting from iteration $k$ of STL ($V_{\phi_{k+1}}$) can directly replace a value model in any search algorithm, such as Greedy Search (§4.1 and §4.2) and BFS (§4.3). As shown in Figure 2 , $V_{\phi_{k+1}}$ simulates a step of lookahead for the state $s_{n}$ i.e. for $({r}_{s_{n}|x},{v}_{s_{n}|x})\sim V_{\phi_{k+1}}(x,s_{0},\ldots,s_{n})$,
@@
where $\tilde{a}_{{n+1}}||\tilde{s}_{{n+1}}$ is a simulated lookahead step and $\tilde{r}_{s_{n+1}|x}$ is its value rationale.
@@ 4 Experiments
We benchmark our proposed STL self-improvement approach on applied web agent tasks, multi-step question answering, and math puzzle tasks [^1] Our code is available at https://github.com/ethanm88/self-taught-lookahead..
4.1 Web Tasks
As mentioned in §1, it is particularly challenging and expensive to gather ground truth web task completion data zhang2024rest_mcts . To benchmark our STL method on web tasks, we utilize WebShop yao2022webshop , which consists of interactive web tasks involving searching for and purchasing an item that matches a short natural language specification. This benchmark is an ideal test bed to demonstrate the ability of our approach, as, unlike other web task datasets zhou2023webarena ; koh2024visualwebarena , ground truth reward is provided for all tasks, allowing a direct comparison between STL and methods in the Reward and Demo Learning and Reward-Guided Inference settings which use this reward.